OpenStudy (anonymous):

Solve the following equation by finding the general solution: 1. 3cosx+1.8=0 2. 2tanx+7tanx+10=0

OpenStudy (matt101):

First rearrange the equation so that each isolates x: FIRST EQUATION: \[3 \cos x+1.8=0\]\[\cos x={1.8 \over 3}\]\[\cos x={3 \over 5}\] SECOND EQUATION: \[2 \tan x+7 \tan x+10=0\]\[9\tan x+10=0\]\[\tan x=-{10 \over 9}\] Note that tanx=sinx/cosx, so we can express the second equation as: \[{\sin x \over \cos x}=-{10 \over 9}\] OR \[\cos x = -{9 \sin x \over 10}\] Now we can set the two equations equal to one another: \[{3 \over 5}= -{9 \sin x \over 10}\] By solving for x in this equation you should be able to get your answer!